# Fidel-TS: A High-Fidelity Benchmark for Multimodal Time Series Forecasting

### v2

@@ Conclusion
Reproducibility Statement
LLM Usage Statement
We used LLMs for grammar checking and rephrasing, without contributing to the core ideas and methodologies.
@@ A.2 Visualization
For illustrative purposes, we selected datasets with a single channel and plotted one of their subjects over a fixed time range. It can be observed from Figure 2 that the CCP and GREG datasets exhibit clear periodicity within the selected range. In contrast, the NYTS dataset is not only noisy but also shows periods of sensor downtime. This suggests that predicting on the NYTS dataset is more challenging and will likely result in higher errors compared to the CCP and GREG datasets.
@@
A.3 Subsetting Method for Generalization Evaluation
@@
Appendix B Framework Comparison
We evaluate our framework against prior ones across several key dimensions: Training Paradigms, Model Type, Modality Support, and Evaluation Strategy. Existing frameworks typically exhibit limitations, lacking comprehensive support for one or more of these dimensions. In contrast, our framework successfully integrates all the aforementioned capabilities (as detailed in Table 7), offering broad and versatile support for model evaluation.
Table 7: Comparison of time series forecasting frameworks across four key dimensions
Benchmarks Training Paradigms Model Type Modality Support Evaluation Strategy
Domain-
specific
Cross-
domain
Nume-
rical
LLM-
based
LLM
API
Uni-
modal
Multi-
modal
Full-shot
Zero-shot
LSTF-Linear
TSLib
TSFM-Bench
$M$-TSFLib
Ours
@@ Appendix C Model Details
Table 8 contains all the models we select for our experiments. We summarize their Training paradigm, Architecture, Modality and Textual Input Requirements. “General" and "channel/subject” corresponds to the dataset descriptions and the meanings of channels/subjects provided in the metadata. “History” denotes the textual information within the historical window. “Condition” refers to the conditions used for forecasting over the future horizon.
@@
PatchTST Domain-Specific Num. trans. Unimodal
iTransformer Domain-Specific Num. trans. Unimodal
Dlinear Domain-Specific Linear Unimodal
FITS Domain-Specific Linear Unimodal
GPT4TS Domain-Specific LLM Unimodal
Chronos Cross-domain Num. trans. Unimodal
TimeMoE Cross-domain Num. trans. Unimodal
Sundial Cross-domain Num. trans. Unimodal
GPT4MTS Domain-Specific LLM Multimodal history
FIATS Domain-Specific Num. trans. Multimodal
condition, general,subject/channel
@@
history, condition, general, subject/channel
@@
history, condition, general, subject/channel
@@
history, condition, general, subject/channel
@@
history, condition, general, subject/channel
@@
condition, general, subject/channel
Figure 4: Prompt template for LLM in unimodal forecasting
Figure 5: Prompt template for LLM in multimodal forecasting
Additionally, to ensure experimental transparency, we provide the prompt templates used for the unimodal and multimodal forecasting tasks with LLMs. The template for unimodal forecasting, illustrated in Figure 4, incorporates only basic information about the dataset and channel, along with the formatted historical time series as input. The system prompt directly instructs the model to learn the pattern of time series and generate the forecast.
The multimodal forecasting template shown in Figure 5 extends the unimodal setup by additionally providing historical textual information and future conditions in the prediction window. The corresponding system prompt directs the LLM to not only learn the pattern of time series but also focus on the association between the textual data and the time series. For both tasks, we explicitly define the output format to minimize the likelihood of formatting errors.
Notably, we design an additional prompt template for the multimodal forecasting of the BEAR dataset to accommodate its unique control events. The weather-related content is replaced with control-related information, while all other parts remained unchanged.
@@ Appendix D Experimental Settings Details
We adopt the Mean Square Error (MSE) as the evaluation metric. The formula is:
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$ (1)
@@ Appendix E Additional Experiment Results
E.1 Benchmarking Unimodal Forecasting without Downsampling
Table 9: Performance of unimodal forecasting models on full test set of each dataset without downsampling. All metrics are MSE.
@@ E.1 Benchmarking Unimodal Forecasting without Downsampling
Patch Dlinear Sundial
@@
We evaluate the unimodal forecasting models on the datasets at their original sampling rates. To maintain the same forecasting setting (a half-month history to predict the next day, the next week, the next 2 weeks and the next month), we adapt the history and prediction lengths to match the native frequency of each dataset. Due to the prohibitive computational cost, we selectively conducted experiments on a subset of models over shorter prediction horizons (day and week), with the results presented in Table 9. These results indicate that the models’ performance on the raw-resolution data is similar to that on the data downsampled to an hourly frequency.
E.2 More Results for Causal Soundness principle
To further investigate the impact of causal relevance on the multimodal prediction performance of FIATS, we report its results on the original, unfiltered full test sets, summarized in Table 10 and 11. The results show that its performance is highly similar to that on the filtered -mini test sets.
A comparison between Table 10 and Table 2 reveals that on datasets with a strong correlation between weather and system dynamics—namely CCP, GREG, JAP, and NYTS—FIATS achieves the best performance among all models on the full test set. Conversely, a comparison of Table 11 and Table 2 shows that for the indoor system represented by the BEAR dataset, using weakly correlated outdoor weather information leads to performance degradation. However, upon switching to highly relevant indoor control events, FIATS once again outperforms all other models. These findings further validate our proposed Causal Soundness principle.
Additionally, we observed a performance decline for FIATS on the CAISO dataset. We are currently unable to conduct a deeper investigation due to the limited variety of available textual information. However, from a causal relevance perspective, we hypothesize that this degradation may be attributed to certain channels in CAISO containing energy information (e.g., biogas, biomass, nuclear) that is weakly associated with weather.
@@ E.2 More Results for Causal Soundness principle
Table 10: Performance of FIATS on the full test set of hourly datasets. All metrics are MSE.
Dataset Pred. Len.
24 168 336 720
CCP-h 0.109 0.159 0.168 0.183
GREG-h 0.044 0.065 0.082 0.094
JAP-h 0.333 0.451 0.500 0.502
CAISO-h 0.202 0.429 0.509 0.600
NYTS-h-obs 0.407 0.644 0.685 0.710
Table 11: Performance of FIATS on the full test set of BEAR datasets. All metrics are MSE.
Textual Input Pred. Len.
12 144 288 576
+Weather 0.129 0.562 0.998 1.309
+Control Event 0.094 0.222 0.298 0.323

### v3

@@ A.2 Visualization
For illustrative purposes, we selected datasets with a single channel and plotted one of their subjects over a fixed time range. It can be observed from Figure 3 that the CCP and GREG datasets exhibit clear periodicity within the selected range. The NYTS dataset is not only noisier and exhibits more frequent, irregular fluctuations, leading to greater prediction difficulty, but also includes periods of sensor downtime, resulting in significant sudden distributional shifts. This suggests that predicting on the NYTS dataset is more challenging and will likely result in higher errors compared to the CCP and GREG datasets.
@@
A.3 Importance Sampling Method for LLM evaluation
Evaluating LLMs on the full test set presents prohibitive computational costs due to the high frequency and large scale of the data. To address this while ensuring the evaluation remains unbiased, we constructed a representative subset (denoted as -mini) using the importance sampling strategy.
We observed that samples within the test set exhibit varying degrees of multimodal dependency: for some instances, textual information is highly correlated with the time series and significantly aids forecasting, while for others, it offers marginal gains. Simple random sampling risks distorting this distribution. To preserve the original "difficulty" distribution, we employed two lightweight models: a representative unimodal model (PatchTST, (Nie et al., 2023)) and a multimodal model (FIATS, (Xu et al., 2024a)). By evaluating both on the full test set, we categorized samples into two groups based on the performance gap: (1) Text-sensitive samples: Instances where the multimodal model significantly outperformed the unimodal one. (2) Text-insensitive samples: Instances where the improvement was negligible.
Finally, we applied stratified sampling with a fixed ratio to each group. Based on the data volume of each dataset, the final selected ratios are: 1% per subject for NYTS, 5% per subject for BEAR, and 10% per subject for other datasets. This ensures an appropriate data volume for the final -mini test set, while test set faithfully reflects the multimodal sensitivity distribution of the original data, providing a reliable and efficient benchmark for LLM evaluation.
A.4 Subsetting Method for Generalization Evaluation
@@
Appendix B Framework Implementation
Our Universal Cross-modal Evaluation Framework supports both raw text and pre-computed embedding vectors while rigorously enforcing the temporal alignment between textual data and time series. To cater to the varied nature of modern models, our framework not only provides standard PyTorch (Paszke et al., 2019) interface, but also integrates HuggingFace Transformers (Wolf et al., 2020) for foundation models, and leverages PyTorch Lightning (Falcon et al., 2019) to accelerate the training of complex multimodal models. Recognizing the unique requirements of LLMs, it also supports both local deployment via vLLM (Kwon et al., 2023) and remote API calls through a simple socket. This thoughtful engineering ensures that our framework serves as the essential bridge between the rich, multimodal data of Fidel-TS and the diverse array of models, enabling the fair and comprehensive experiments that follow.
@@ Appendix C Model Details
Table 7 contains all the models we select for our experiments. We summarize their Training paradigm, Architecture, Modality and Textual Input Requirements. “General Info" and "Channel/Subject Info” corresponds to the dataset descriptions and the meanings of channels/subjects provided in the metadata. “History” denotes the textual information within the historical window. “Condition” refers to the conditions used for forecasting over the future horizon.
@@
PatchTST Domain-Specific Numerical Transformer Unimodal
iTransformer Domain-Specific Numerical Transformer Unimodal
Dlinear Domain-Specific Linear (MLP) Unimodal
FITS Domain-Specific Linear (MLP) Unimodal
GPT4TS Domain-Specific LLM-Reprogramming Unimodal
Chronos Cross-Domain Numerical Transformer Unimodal
TimeMoE Cross-Domain Numerical Transformer Unimodal
Sundial Cross-Domain Numerical Transformer Unimodal
GPT4MTS Domain-Specific LLM-Reprogramming Multimodal History
FIATS Domain-Specific Numerical Transformer Multimodal
Condition, General Info, Subject/Channel Info
@@
History, Condition, General Info, Subject/Channel Info
@@
History, Condition, General Info, Subject/Channel Info
@@
History, Condition, General Info, Subject/Channel Info
@@
History, Condition, General Info, Subject/Channel Info
@@
Condition, General Info, Subject/Channel Info
@@ Appendix E Additional Experiment Results
E.1 Results of Unimodal Forecasting Models on Datasets without Downsampling
To maintain the same forecasting setting (a half-month history to predict the next day, the next week, the next 2 weeks and the next month), we adapt the history and prediction lengths to match the native frequency of each dataset. Due to the prohibitive computational cost, we selectively conducted experiments on a subset of models (for transformer-based, MLP-based, and foundation models, we select one representative model from each category) over shorter prediction horizons (day and week), with the results presented in Table 8. These results indicate that the models’ performance on the raw-resolution data is similar to that on the data downsampled to an hourly frequency.
Table 8: Performance of selected unimodal forecasting models on full test set of each dataset without downsampling. All metrics are MSE.
Dataset
@@ E.1 Results of Unimodal Forecasting Models on Datasets without Downsampling
Patch. Dlinear Sundial
@@
E.2 Generalization Test Results of BEAR Dataset
The generalization results for the BEAR dataset using our proposed strategy are presented in Table 9. Consistent with the findings on NYTS, the model rankings in this setting align with those obtained from the in-domain evaluation.
Table 9: New generalization test of domain-specific unimodal forecasting models. All metrics are MSE. The best results are highlighted in bold.
Dataset Pred. Len. Patch. iTrans. Dlinear FITS GPT4TS
BEAR-obs $\rightarrow$ BEAR-hid 12 0.100 0.059 0.105 0.139 0.094
144 0.424 0.382 0.456 0.507 0.426
288 0.588 0.549 0.595 0.659 0.579
576 0.762 0.717 0.801 0.906 0.876
E.3 Evaluation Results of LLMs on Multivariate Datasets
In multivariate forecasting scenarios, the simultaneous processing of numerous variables results in excessive token lengths, frequently causing standard 128k-context LLMs to fail due to context overflow. Consequently, we restrict our evaluation on the multivariate -mini test set to models capable of supporting extended context windows, with the performance results summarized in Table 10. Despite this filtration, we observe two distinct failure modes in these tasks. First, even when operating within theoretical limits, models often struggled to generate the full sequence, resulting in truncated outputs. Second, the overwhelming context length frequently impaired instruction adherence. Models tended to overlook the specific constraint to "predict multivariate variables while maintaining dimensions," resulting in outputs with incorrect feature dimensionality rather than the required original structure.
Table 10: Performance of LLMs on -mini test set of multivariate dataset. All metrics are MSE. Considering the limitations of LLM context, we choose the shorter prediction length (24, 168, and for BEAR is (12, 144)). Blank indicates that the filled-in prompt exceeds the input length limit of the LLM.
Model / Setting JAP -h-mini CAISO -h-mini BEAR -mini
24 168 24 168 12 144
Qwen2.5- 14B-Inst. 1M Uni. 1.825 0.501 0.522 7.425
Mul. 1.457 0.421 1.577(+Wea.) – (+Ctrl.)
Chattime Uni. 0.483 0.580 0.216 0.453 0.145 1.077
Mul. 0.624 0.874 1.209(+Wea.) – (+Ctrl)
E.4 Full Results of Forecasting Models on -mini Test Sets
To facilitate a direct comparison between LLMs and specialized forecasting models, we evaluate all baseline models on the -mini test sets, as summarized in Table 11. The performance trends and model characteristics observed on this subset align closely with those from the full test sets. This consistency attests to the fairness of our sampling algorithm, confirming that the -mini test sets effectively reduce the number of samples and thereby lower the computational overhead for LLM evaluation, while maintaining representativeness and avoiding the introduction of model-specific biases.
@@ E.4 Full Results of Forecasting Models on -mini Test Sets
Table 11: Performance of all forecasting models on -mini test set of all datasets. All metrics are MSE.
Dataset Pred. Len. Domain-Specific (Unimodal) Foundation Models (Unimodal) Domain-Specific (Multimodal)
PatchTST iTrans. Dlinear FITS GPT4TS Chronos Time-MoE Sundial GPT4MTS FIATS
CCP-mini 24 0.140 0.127 0.138 0.141 0.129 0.144 0.146 0.131 0.146 0.093
168 0.192 0.169 0.197 0.197 0.179 0.203 0.219 0.192 0.199 0.142
GREG-mini 24 0.073 0.070 0.089 0.086 0.090 0.093 0.089 0.085 0.092 0.047
168 0.135 0.135 0.150 0.137 0.143 0.170 0.162 0.154 0.156 0.074
JAP-mini 24 0.286 0.285 0.334 0.334 0.278 0.277 0.278 0.286 0.275 0.234
168 0.418 0.411 0.419 0.429 0.421 0.443 0.449 0.448 0.416 0.304
NYTS-mini 24 0.826 0.764 0.864 0.871 0.774 0.880 0.923 0.777 0.838 0.387
168 1.538 1.599 1.560 1.615 1.749 1.876 1.675 1.797 1.643 0.700
CAISO-mini 24 0.136 0.140 0.143 0.145 0.137 0.146 0.132 0.130 0.138 0.170
168 0.276 0.297 0.324 0.285 0.278 0.297 0.298 0.282 0.303 0.364
BEAR-mini 12 0.103 0.080 0.111 0.144 0.091 0.126 0.118 0.116 0.093(Ctrl.) 0.102(Wea.) 0.066(Ctrl.) 0.123(Wea.)
144 0.484 0.353 0.505 0.593 0.432 0.836 0.831 1.027 0.540(Ctrl.) 0.569(Wea.) 0.211(Ctrl.) 0.546(Wea.)
E.5 Full Results of Pass@3 in LLM evaluation
To systematically evaluate the reliability of LLMs in forecasting tasks, we report the pass@3 metric (Table 12), which reveals critical stability bottlenecks tied to input complexity and context limitations. The results highlight three primary failure mechanisms:
First, the impact of High-Dimensional Data and Context Explosion. In unimodal settings, LLMs exhibit high stability on datasets with fewer variables (e.g., CCP, GREG, NYTS), maintaining pass rates above 90%. However, performance collapses on high-dimensional multivariate datasets like JAP and CAISO. Standard 128K-context models yield a 0 pass rate on these tasks. This is attributed to context explosion, where the sheer volume of historical numerical tokens from multiple channels exceeds the model’s effective window, preventing valid output generation.
Second, the "Multimodal Tax" on Reliability. Introducing textual information significantly degrades stability compared to unimodal baselines. Unlike prior works using sparse text, our benchmark involves high-frequency, dense textual data. This additional token load exacerbates the context bottleneck. For instance, on the BEAR dataset, even robust models like DeepSeek-R1 drop from a 99.9% pass rate (Unimodal) to 0 (Multimodal) as the combined input length of weather reports and historical data overwhelms the context window. Similarly, the domain-finetuned Chattime, while achieving perfect stability (100%) in unimodal tasks, fails catastrophically in multimodal long-horizon scenarios (dropping to 0.0%), indicating that current fine-tuning strategies may not sufficiently adapt models to handle such dense multimodal contexts.
Third, the prediction Horizon Sensitivity. Across all datasets and models, extending the prediction horizon (e.g., from 24 to 168) consistently lowers pass rates. This suggests that LLMs struggle to maintain coherence over long output sequences, often resulting in truncation or format deviations when generating extended forecasts. In summary, while LLMs show promise in low-resource settings, their reliability is severely compromised in realistic scenarios requiring the simultaneous processing of high-dimensional variables, long-term history, and extensive textual context.
Table 12: Summarization of all pass@3 between Unimodal (Uni.) and Multimodal (Multi.) forecasting for LLMs.
Dataset Pred. Len. Qwen2.5-14B-Instruct Qwen3-14B Chattime DeepSeek-R1
128K 1M 128K 128K
Uni. Mul. Uni. Mul. Uni. Mul. Uni. Mul. Uni. Mul.
CCP-mini 24 96.3 96.8 96.9 97.1 98.8 95.6 100.0 100.0 99.9 99.8
168 75.0 71.4 73.6 81.5 90.6 77.2 100.0 0.0 97.5 94.9
GREG-mini 24 96.9 98.1 97.6 99.4 96.7 94.3 100.0 99.2 99.8 99.8
168 74.4 91.6 73.2 86.4 88.0 77.2 100.0 0.0 93.2 97.2
JAP-mini 24 0.0 0.0 29.1 47.4 0.0 0.0 100.0 100.0 0.0 0.0
168 0.0 0.0 0.0 0.0 0.0 0.0 100.0 0.0 0.0 0.0
NYTS-mini 24 93.2 95.7 93.7 97.1 97.7 96.3 100.0 99.9 99.9 100.0
168 72.9 81.8 69.4 82.4 90.0 90.0 100.0 0.0 95.3 98.2
CAISO-mini 24 0.0 0.0 45.8 69.6 0.0 0.0 100.0 48.9 0.0 0.0
168 0.0 0.0 0.0 0.0 0.0 0.0 100.0 0.0 0.0 0.0
BEAR-mini 12 96.7 0.0 95.2 63.6(+Wea.) 0.0(+Ctrl.) 99.8 0.0 100.0 88.3(+Wea.) 0.0(+Ctrl.) 99.9 0.0
144 11.6 0.0 58.1 0.0 62.8 0.0 100.0 0.0 88.4 0.0
Appendix F LLM Evaluation Details
F.1 LLM Prompt Template
We provide the prompt templates used for the unimodal and multimodal forecasting tasks with LLMs.
The template for unimodal forecasting, illustrated in Figure 5 (a), incorporates only basic information about the dataset and channel, along with the formatted historical time series as input. The system prompt directly instructs the model to learn the pattern of time series and generate the forecast.
The multimodal forecasting template shown in Figure 5 (b) extends the unimodal setup by additionally providing historical textual information and future conditions in the prediction window. The corresponding system prompt directs the LLM to not only learn the pattern of time series but also focus on the association between the textual data and the time series. For both tasks, we explicitly define the output format to minimize the likelihood of formatting errors.
Notably, we design an additional prompt template for the multimodal forecasting of the BEAR dataset to accommodate its unique control events. The weather-related content is replaced with control-related information, while all other parts remained unchanged. We applied these templates only to native LLMs, including the Qwen series and Deepseek-R1. For the Chattime which fine-tuned Llama-2-7b, we utilized its original template.
Figure 5: Prompt template for LLM unimodal forecasting and multimodal forecasting
F.2 Case Study: Example of LLM Reasoning on a Forecasting Task
To illustrate how LLMs utilize known information within the context in unimodal and multimodal forecasting scenarios, we compared the reasoning processes of Deepseek-R1 under both tasks. The examples we use are derived from the CCP-mini test set. The results are shown in Figure 6.
In the unimodal forecasting scenario, the model relies exclusively on historical time-series data and implicit domain knowledge. The reasoning process focuses on identifying periodicity (daily solar cycles) and seasonal trends (increasing daylight in the Northern Hemisphere). Crucially, the model explicitly acknowledges the lack of external variables, stating, "I don’t have weather data." Consequently, it defaults to a statistical approach, assuming that recent weather will persist, and generates predictions by applying numerical multipliers derived solely from historical correlations.
In contrast, the multimodal forecasting scenario demonstrates a more robust reasoning capability by integrating textual weather data with numerical history. The model actively utilizes the provided weather forecast ("partly sunny skies") to select a specific reference day (March 10) that shares similar conditions. Instead of simple extrapolation, it employs a logic-driven adjustment: it calculates a ratio based on the difference between the current day and the reference day at a known timestamp (11 AM) and applies this scaling factor to forecast the subsequent hours. This highlights the model’s ability to ground its predictions in semantic context rather than relying solely on historical data.
Figure 6: Comparison of the Reasoning Processes of Deepseek-R1 on Unimodal and Multimodal Forecasting Tasks